Accounts Receivable Analysis The following data are taken from the financial statements of McKee Technology Inc. Terms of all sales are 2/10, n/55. 2014 2016 $212,000 1,335,900 1,284,800 2015 Accounts receivable, end of year Net sales on account a. For 2015 and 2016, determine (1) the accounts recelvable turnover and (2) the number of days sales in receivables. Round answers to one decimal place. Assume a 365-day year. $226,000 $241,200 2016 2015 1. Accounts receivable turnover 2. Number of days sales in receivables b. The collection of accounts receivable has improved days days . This can be seen in both the increase in accounts receivable turnover and the decrease in the collection period
Answer #1

Calculate following :

2016 2015
Average receivable (212000+226000/2) = 219000 (226000+241200/2) = 233600
1. Account receivable turnover 1335900/219000 = 6.1 1284800/233600 = 5.5
2. Number of days sales in receivable 365/6.1 = 59.8 Days 365/5.5 = 66.4 days

b) The collection of account receivable has improved. This can be seen in both the increase in account receivable turnover and the decrease in collection period